OpenStudy (anonymous):

A screening test for a disease shows a positive test result in 95% of all cases when the disease is actually present and in 15% of all cases when it is not. When the test was administered, 17.4% of the results were positive. What is the prevalence of the disease?

OpenStudy (kirbykirby):

Let \(T\) = test is positive \(D \) = disease present \(\overline{D}\)= disease absent You are given: \(P(T|D)=0.95\\ P(T|\overline{D})=0.15 \\ P(T)=0.174\) You are asked: to find \(P(D)\), which is the prelevance of the disease So: \(P(T)=P(T \cap D) +P(T \cap \overline{D})\), by law of total probability since \(D \) and \(\overline{D}\) form a partition \(P(T)=P(T|D)P(D) +P(T|\overline{D})P(\overline{D})\\ 0.174=0.95(P(D))+0.15(1-P(D))\\ 0.174=0.95P(D)+0.15 - 0.15P(D)\\ 0.024 =0.8P(D)\\ P(D)=0.024/0.8=0.03=3\%\)